In a radial electrostatic filter, find the field strength in Volts per centimeter, necessary to bend

singly ionized atomic Oxygen ions in an arc having a 25 cm radius if the ions have fallen through

a potential difference of 20 000 V before entering the filter.


1
Expert's answer
2021-03-16T08:29:49-0400
"E=\frac{U}{d}\\E=\frac{20000}{25(2)}=400\frac{V}{cm}"